Determine the amount of sales (units) that would be necessary under Break-Even Sales Under Present and Proposed Conditions Darby Company, operating at full capacity, sold 120,150 units at a price of $114 per unit during the current year. Its income statement for the current year is as follows: Sales $13,697,100 Cost of goods sold 6,764,000 Gross profit $6,933,100 Expenses: Selling expenses $3,382,000 Administrative expenses 3,382,000 Total expenses 6,764,000 Income from operations $169,100 The division of costs between fixed and variable is as follows: Variable Fixed Cost of goods sold 70% 30% Selling expenses 75% 25% Administrative expenses 50% 50% Management is considering a plant expansion program that will permit an increase of $1,254,000 in yearly sales. The expansion will increase fixed costs by $125,400, but will not affect the relationship between sales and variable costs. Required: 1. Determine the total variable costs and the total fixed costs for the current year. Enter the final answers rounded to the nearest dollar. Total variable costs $ Total fixed costs $ 2. Determine (a) the unit variable cost and (b) the unit contribution margin for the current year. Enter the final answers rounded to two decimal places. Unit variable cost $ Unit contribution margin $ 3. Compute the break-even sales (units) for the current year. Enter the final answers rounded to the nearest whole number. units 4. Compute the break-even sales (units) under the proposed program for the following year. Enter the final answers rounded to the nearest whole number. units 5. Determine the amount of sales (units) that would be necessary under the proposed program to realize the $169,100 of income from operations that was earned in the current year. Enter the final answers rounded to the nearest whole number. units 6. Determine the maximum income from operations possible with the expanded plant. Enter the final answer rounded to the nearest dollar. $ 7. If the proposal is accepted and sales remain at the current level, what will the income or loss from operations be for the following year? Enter the final answer rounded to the nearest dollar. $ 8. Based on the data given, would you recommend accepting the proposal? In favor of the proposal because of the reduction in break-even point. In favor of the proposal because of the possibility of increasing income from operations. In favor of the proposal because of the increase in break-even point. Reject the proposal because if future sales remain at the current level, the income from operations will increase. Reject the proposal because the sales necessary to maintain the current income from operations would be below the current year sales.
Answer #1

Question 1

Cost of Goods Sold = $ 67,64,000

Variable Costs of Goods Sold = 67,64,000 * 70% = $ 47,34,800

Fixed Costs of Goods Sold = 67,64,000 * 30% = $ 20,29,200

Selling Expense = $ 33,82,000

Variable Selling Expenses = 75% of 33,82,000 = $ 25,36,500

Fixed Selling Expenses = 25% of 33,82,000 = $ 845,500

Administrative Expenses = $ 33,82,000

Variable Administrative Expenses = 50% of 33,82,000 = $ 16,91,000

Fixed Admisnistrative Expense = 50% of 33,82,000 = $ 16,91,000

Total Variable Expenses = Variable Cost of Goods Sold + Variable Selling Expenses + Variable Administrative Expenses

Total Variable Expenses = 47,34,800 + 25,36,500 + 16,91,000

Total Variable Costs = $ 89,62,300

Total Fixed Expenses = Fixed Cost of Goods Sold + Fixed Selling Expenses + Fixed Administratives Expenses

Total Fixed Expenses = 20,29,200 + 8,45,500 + 16,91,000

Total Fixed Costs = $ 45,65,700

Question 2

Variable Costs per Unit = Total Variable Costs / Total Units

Variable Costs per Unit = 89,62,300 / 120,150

Variable Costs per Unit = $ 74.59

Contribution Margin per Unit = Sales Price per Unit - Variable Costs per Unit

Contribution Margin per Unit = 114 - 74.59

Contribution Margin per Unit = $ 39.41

Question 3

Break Even Point in Units = Total Fixed Costs / Contribution Margin per Unit

Break Even in Units = 45,65,700/ 39.41

Break Even Point in Units = 115,851 Units

Question 4

Break Even Point in Units = Total Fixed Costs / Contribution Margin per Unit

Total Fixed Costs = 45,65,700 + 125,400(Additional Fixed Costs )

Total Fixed Costs = $ 46,91,100

Break Even Point in Units = 46,91,100 / 39.41

Break Even Point in Units = 119,033 Units

Question 5

Contribution Margin to earn Operating Income of $ 169,100 in proposed program = Fixed Cost Propwith Newosed + Operating Income of 169,100

Fixed Costs with New Proposed Program = $ 46,91,100

Contribution Margin to earn Operating Income of $ 169,100 = 46,91,100 + 169,100 = $ 48,60,200

Units to be sold to earn Contribution Margin of $ 48,60,200 = 48,60,200 / Contribution Margin per Unit

Unit to be Sold ro earn Contribution Margin of $ 48,60,200 = 48,60,200 / 39.41 =

= 123,324 Units

To Earn $ 169,100 under Proposed Program 123,324 Units are required to be sold.

Question 6

Maximum Income from Operation from Proposed Program = Total Sales - Total Variable Costs - Total Fixed Costs

Total Fixed Costs = 46,91,100

Total Sales = 1,36,97,100 + 12,54,000 (Additional Sales ) = $ 1,49,51,100

Total Variable Costs = 74.59 per Unit * 131,15000 Units = $ 97,82,479

Maximum Income from Operation = 1,49,51,100 - 97,82,479 - 46,91,100

Maximum Income from Operation =$ 477,521

Note

Total Units = 120,150 + 11,000 Unit = 131,150 Units

ADDITIONAL Unit's = 12,54,000 / $ 114 per Unit = 11,000 Units

Question 7

No Increase in Sale and Variable Costs

Income / (Loss) from Operation = Total Sales - Total Variable Costs - Total Fixed Costs

Total Fixed Costs = 46,91,100

Total Sales = 1,36,97,100

Total Variable Costs = $ 89,62,300

Income / (Loss) from Operation = 1,36,97,100 - 89,62,300 - 46,91,100

Income / (Loss) from Operation = $ 43,700

Question 8

In favour of the proposal because of the possibility of increasing Income from Operations.